OpenStudy (anonymous):

Help with Linear Algebra Question Please. Find the cubic polynomial f(x) such that f(2)=-13, f'(2)=-11, f''(2)=-10, and f'''(2)=-6. f(x)= ___ x^3+ ____x^2+____ x+ ____. I know to find the integral and i found these (2,-13) f(2) = -11x f(2) = -5x^2 f(2) = -x^3 But i'm confused of what to do next. Thanks a lot

OpenStudy (anonymous):

\(f(x)=ax^3+bx^2+cx+d\) \(f'(x)=3ax^2+2bx+c\) \(f''(x)=6ax+2b\) \(f'''(x)=6a\) now change x by 2 and solove system for a,b,c,d

OpenStudy (anonymous):

so 8a + 4b + 2c + d = -13 12a + 4b + c = -22 12a + 2b = -20 6a=-8?

OpenStudy (anonymous):

well i know a is -1 cause i guessed :3 but it seems like it's wrong in this case

OpenStudy (anonymous):

f'''(2)=-6. so 6a=-6 a=-1

OpenStudy (anonymous):

it's right

OpenStudy (anonymous):

how did you get -6?

OpenStudy (anonymous):

it's written in your problem

OpenStudy (anonymous):

"Find the cubic polynomial f(x) such that f(2)=-13, f'(2)=-11, f''(2)=-10, and f'''(2)=-6. "

OpenStudy (anonymous):

aaaah okay. so finding the integral doesn't have to do with this

OpenStudy (anonymous):

so the system will be: 8a+4b+2c+d=-13 12a+4b+c=-11 12a+2b=-10 6a=-6 you don't need integral

OpenStudy (anonymous):

ok. i get it. thanks a lot!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so a = -1 b = 1 c = -3 and d = -3 and they're right :) thank you again!

OpenStudy (anonymous):

yw